# arXiv:2603.03220v1[cond-mat.mtrl-sci] 03 Mar 2026

### v1

@@ Abstract
A novel implementation of the linear response time-dependent density functional theory addressing spin excitations in non-collinear magnets based on the Korringa-Kohn-Rostoker Green’s function method is presented. Following the exposition of the formalism based on the adiabatic local spin density approximation to the exchange-correlation kernel generalized to the non-collinear case, the computational scheme is discussed in detail. The formation of the Goldstone modes in non-collinear susceptibility calculations is elaborated on formally and from the numerical convergence point of view. The scheme is deployed to study the dispersion, Landau damping, and spatial shapes of magnons for the representative members of the kagome non-collinear antiferromagnets.
@@ 1 Introduction
The ab initio linear response time-dependent density functional theory (LRTDDFT) of spin excitations has become an indispensable tool in the computational design of novel functional materials [Gross1985, Savrasov1998, Staunton1999, Buczek2009, Durhuus2023, Binci2025]. It offers a comprehensive description not only of spin-waves (magnons), which can be also studied by means of an empirical mapping onto the celebrated Heisenberg model [LeBlanc2021], but also of Stoner excitations and of the associated Landau damping, inherent to the itinerant magnetism [Moriya1985]. No doubt the dire need of the moment is the extension of the LRTDDFT to non-collinear (NC) magnets, i.e., spin-polarized solids where the direction of magnetization is allowed to vary as a function of position [Sandratskii1998]. In the spintronic and magnonic context, an enormous practical potential has been ascertained to the members of this material family [1, Rimmler2024]. Moreover, they often accompany other exotic or strongly correlated phases of matter, e.g., the superconductivity [Bruening2025]. However, the true attractive force of these magnets stems from their beautiful symmetries [Sandratskii1991] rooted deeply in the modern field theories [Azaria1993, Pradenas2024, Pradenas2025]
@@
A pivotal role in the development of magnonic devices is played by the damping of the spin-waves [Chumak2015, Joshi2016, Duine2018]. Except for the Landau mechanism, spin-waves can also scatter considerably on crystal imperfections [Buczek2016]. Remarkably, NC magnets cannot be fully grasped within the linear spin-wave theory and their magnons undergo intrinsic damping due to two-boson interactions even at absolute zero [Chernyshev2009, Zhitomirsky2013, Park2020] contrary to the case of collinear ferromagnets [Paischer2024].
Prompted to close this important methodological gap, in this paper, we present the first computer implementation of the LRTDDFT for NC magnets and deploy it to study collective spin excitations in representative members of the NC kagome antiferromagnets (KAF) [Park2018, Park2020].
@@
The paper is organized as follows. Sections 2 and 3 expose the underlying formalism and our implementation based on the Korringa-Kohn-Rostoker Green’s function method. In section 4, we discuss the dispersion, Landau damping, and spatial shapes of magnons for the representative members of the NC KAF.
@@ 2.1 Generalized susceptibility
In the linear regime, the quantity yields the charge or magnetization density response $\delta n^{i}{\left({\mathbf{x}},t\right)}$ emerging due to the dynamical perturbing fields $\delta V^{j}{\left({\mathbf{x}}^{\prime},t^{\prime}\right)}$ (magnetic or scalar) coupling to these densities. $\hat{\sigma}^{i}{\left({\mathbf{x}}t\right)}$ are charge ($i=0$) and magnetization density operators ($i=x,y,z$), ${{\left[A,B\right]}}\equiv AB-BA$ is the commutator, and the $^{\left\langle\hat{o}\right\rangle}$ is the expectation value of the operator $\hat{o}$ for the system without the perturbation. Throughout all the paper, $\mathbf{x}$} refers to the global position vector, whereas $\mathbf{r}$} is the position within the primitive unit cell. It is assumed that the potentials of the unperturbed system bear no time dependence in the Schrödinger picture. Note that the response function can be used to study any electron density excitations including magnons, excitons, and plasmons.
@@
Here, $\mathbf{r}$} and $\mathbf{r}$$^{\prime}$ belong to the same primitive cell of the crystal and $\mathcal{L}$ denotes the set of lattice vectors. (In the equation above other arguments are suppressed for simplicity.)
@@
allows to analyse the spectra and the spatial forms of the excited density modes. For a specific energy $\omega$, the eigenvectors $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ of $L$ represent the shapes of the natural modes of charge and magnetization density fluctuations at this frequency. The magnitude of the associated eigenvalues, $a_{\lambda}{\left({\mathbf{q}},\omega\right)}$, gives the density of states of these modes, up to a factor $1/\pi$. In other words, large $a_{\lambda}$ signifies the presence of collective density modes of the system. These modes are real physical objects of the following form
${\mathfrak{Re}{\left[\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}{\mathfrak{e}^{{{\mathfrak{i}}}{\left({\mathbf{q}}\cdot{\mathbf{R}}-\omega t-\phi\right)}}}\right]}},$ (4)
where $\phi$ is an overall phase of the mode. Following the fluctuation dissipation theorem, which we briefly outline in Appendix A, the loss matrix has several further useful physical interpretations. On one hand, $\omega a_{\lambda}{\left({\mathbf{q}},\omega\right)}/2$ yields the power absorbed by the system from the driving external fields of the shape $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ coupling to the densities. On the other, it can be used to compute the expectation values of the time correlations ${\mathfrak{C}}^{ij}{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}},\omega\right)}$ between density operators in the unperturbed state, i.e., among others, the spontaneous fluctuations of the system density.
Furthermore, the magnetization density loss matrix evaluated in the momentum space is proportional to the cross-section of the inelastic neutron scattering off spin-density excitation [VanHove1954]. Last but not least, the quantity is related to the (spin-polarized) electron energy loss spectroscopy ((SP)EELS) signal [Kirschner1985, Vollmer2003, Costa2006, Santos2018]. In the latter case, however, contrary to the neutron scattering, the loss energy dependence of the cross-section involves intricate details of the Coulomb interaction between the electrons of the beam and the probe [Vignale1985, Plihal1998, Hong2000a, Paischer2024b].
@@ 2.2 Linear response time dependent density function theory formulation
The evaluation of the dynamic susceptibilities of interacting electron gas is as involved as the solution of the many-body problem itself [Abrikosov1975] and one must resort to approximations. A successful ab initio scheme can be derived within the LRTDDFT [Gross1985, Qian2002] and it often involves two steps. First, the so called Kohn-Sham susceptibility, $\chi_{\mathrm{KS}}$ is constructed describing the response of the formally non-interacting Kohn-Sham (KS) system to an external perturbation. Second, the self-consistent dynamical change in the exchange-correlation potential arising due to the densities induced by the external field and this change is incorporated upon solving so called susceptibility Dyson equation. There are different complementary approaches to implement this scheme and some do not explicitly break it into these two distinct steps [Savrasov1998, Binci2025]. The magnetic response function can be also found within the many-body perturbation scheme [Aryasetiawan1999, Karlsson2000, Mueller2016, Nabok2021] and using empirical Hubbard-type Hamiltonians [Cooke1980, Gumbs1980, Tang1998, Muniz2002, Bonetti2022].
@@
(For simplicity the arguments of the Green’s functions as well as corresponding sums and integrals are suppressed.) The ground state potential $V$ includes the static external fields as well as the ground state Hartree and exchange-correlation contributions. In this case it does not feature the dependence on the complex energy $z$ and is local. This is the form used to describe the materials considered in this paper.
Additionally, in many cases, it is necessary to include terms originating from the spin-orbit coupling (SOC) [Paischer2024]. Furthermore, the frequency dependence arises when typical electronic self-energy terms are incorporated, as it is the case when the disorder is described in the coherent potential approximation (CPA) [Buczek2016, Paischer2021, Paischer2021a] or many-body effects like magnon-electron interaction are included [Paischer2023, Usachov2024, Paischer2024a]. In the latter case the self-energy $V$ can additionally become non-local.
@@
${f_{T}}({\epsilon}_{j})$ stands for the Fermi-Dirac distribution function. A small finite temperature is introduced to smear slightly the Fermi level allowing for better convergence of the Brillouin zone integrals. In this work, this temperature is negligible on the electronic energy scales and the electron gas is regarded to be described at absolute zero. $0^{+}$} denotes the small positive energy limit. Einstein summation convention is assumed for the spin indices and there is no difference raised and lowered density index $i$. The advantage of this approach is the possibility of deforming the complex plane integration contours such that the evaluation of $S$ takes place far from the singularities of the KS Green’s function allowing for its more rapid numerical convergence [Buczek2011a]. Furthermore, the evaluation of the susceptibility on the nearly real axis yields more stable analytic continuation to the real frequencies than the evaluation at Matsubara poles [Staunton1999, Staunton2000, Thakor2003].
@@ 2.3 Exchange-correlation kernel
The exact form of the $K_{\mathrm{xc}}$ is, in general, unknown, especially for NC systems. There is a steady progress in constructing NC spin density functionals [Sharma2007, Kurth2009, Eich2010, Eich2013, Eich2013a, Pittalis2017, Eich2018], but their their inclusion in practical LRTDDFT schemes is still in progress. In this work, we adopt the adiabatic local spin density approximation (ALSDA) yielding $K_{\mathrm{xc}}$ in the following form
@@
For NC systems, this kernel is most easily written in a local coordinate system with its $z$-axis oriented along the magnetization direction [vonBarth1972, Katsnelson2004]. Its space-diagonal part reads in this case
$K_{\mathrm{xc}}^{\textrm{loc}}=\begin{pmatrix}{\frac{\partial V_{\mathrm{xc}}}{\partial m}}&0&0&{\frac{\partial V_{\mathrm{xc}}}{\partial m}}\\ 0&K^{\perp}&0&0\\ 0&0&K^{\perp}&0\\ -{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial n}}&0&0&-{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial m}}\\ \end{pmatrix}$ (12)
@@
$K^{\perp}{\left({\mathbf{x}}\right)}$ $=-{\mu_{\textrm{B}}}\frac{B^{\mathrm{GS}}_{\gamma}{\left({\mathbf{x}}\right)}}{m^{\mathrm{GS}}_{\gamma}{\left({\mathbf{x}}\right)}}$ (14)
$K^{\perp}$ is found simply as the ratio of the magnitudes of KS magnetic field and magnetization density. We recall that these two latter vectors are parallel in the LSDA. Furthermore, it is immediately evident that $K^{\perp}$ is isotropic in the subspace of the transverse magnetization direction. This form has a clear physical interpretation. Once a infinitesimal magnetization is induced in the direction perpendicular to ${\mathbf{m}}^{\mathrm{GS}}{\left({\mathbf{x}}\right)}$ the local ${\mathbf{B}}_{\mathrm{xc}}{\left({\mathbf{x}}\right)}$ rotates locally without changing its magnitude.
@@
2.4 The emergence of the Goldstone modes in non-collinear magnets
@@
While the emergence of the three Goldstone bosons in NC triangular ferromagnets can be be deduced from the general symmetry arguments outlined above, it is instructive to inquire how it manifests itself in the LRTDDFT calculations. Here, the correct long-wavelength behaviour, i.e., the presence of a mode or modes with vanishing energy for $q=0$, can be seen as a sum rule arising from the interplay between the exchange correlation kernel $K_{\mathrm{xc}}$ and the static uniform magnetic susceptibility [Lounis2010, Lounis2011, Buczek2011a, Mueller2016]. We name this interplay Goldstone sum rule and it can be stated as follows. The “Dyson denominator” of Eq. (2.2) for $q=0$ and $\omega=0$ (Coulomb interaction, $v_{\textrm{C}}$ plays no role in the formation of the magnetic NGB with anisotropy neglected)
@@
must feature zero eigenvalues. Their multiplicity, $N_{r}$, is equal to the number of linearly independent rigid rotations of the magnetic order parameter which, with spin-orbit coupling neglected, cost no energy. We refer to them as zero-energy magnetic deformations. They are the eigenvectors of the Dyson denominator with eigenvalue 0. In a NC magnets $N_{r}=3$ while $N_{r}=2$ holds for the collinear case, i.e., it is directly given by the number of broken symmetry generators. However, in the light of the preceding discussion, we note that $N_{r}$ is not necessarily equal to the number of Goldstone modes of a system.
Let us turn now to our specific case of the ALSDA. Apart from having no frequency dependence, this kernel, cf. Sec. 2.3, is $q$-independent and finite at $q=0$. The latter property tends to be a serious drawback in the charge-charge response channel, as it precludes the correct TDDFT description of excitonic effects in semiconductors [Onida2002, Turkowski2009]. On the other hand, it is the correct behaviour in the long wave-length limit for magnetization density channels, yielding correctly the magnon Goldstone mode which would not arise otherwise.
@@
The first is of systematic character and arises often in the many-body perturbation methods [Mueller2016, Nabok2021] where the ground state found self-consistently using an approach (typically LSDA to the ground state functional) different from the one for the kernel (corresponding to the screened Coulomb interaction $W$). Once $W$ is used to self-consistently find the magnetic ground state the deficiency will disappear. Let us briefly remark that the currently available many-body perturbation schemes for computing magnetic excitations are practically equivalent to the LRTDDFT in the ALSDA. They take advantage of the LSDA for the electronic ground state, local and static approximation to the screened Coulomb interaction $W$, as well as the choice of Feynman diagrams yielding the random phase approximation-like equation for the dynamic susceptibility resembling formally the equation (2.2) of the LRTDDFT.

### v2

# Dynamical magnetic susceptibility of non-collinear magnets: a novel KKR-based ab initio scheme and its application

@@ Abstract
A novel implementation of the linear response time-dependent density functional theory addressing spin excitations in non-collinear magnets based on the Korringa-Kohn-Rostoker Green’s function method is presented. Following the exposition of the formalism based on the adiabatic local spin density approximation to the exchange-correlation kernel generalized to the noncollinear case, the computational scheme is discussed in detail. The formation of the Goldstone modes in non-collinear susceptibility calculations is elaborated on formally and from the numerical convergence point of view. The scheme is deployed to study the dispersion and Landau damping of magnons in the altermagnetic non-collinear kagome antiferromagnet Mn3Ir. The non-monotonous dependence of the damping on the magnon frequency makes the large momentum excitations attractive in the terahertz spintronics. To this end, we analyze the real-time and real-space dynamics of the magnetic modes, including their strongly chirality-dependent attenuation.
@@ 1 Introduction
The ab initio linear response time-dependent density functional theory (LRTDDFT) of spin excitations has become an indispensable tool in the computational design of novel functional materials [Gross1985, Savrasov1998, Staunton1999, Buczek2009, Lounis2011, Skovhus2021, Skovhus2022, Durhuus2023, Binci2025]. It offers a comprehensive description not only of spin-waves (magnons), which can be also studied by means of an empirical mapping onto the celebrated Heisenberg model [LeBlanc2021], but also of Stoner excitations and of the associated Landau damping, inherent to the itinerant magnetism [Moriya1985]. No doubt, the dire need of the moment is the extension of the LRTDDFT to non-collinear (NC) magnets, i.e., spin-polarized solids where the direction of magnetization is allowed to vary as a function of position [Sandratskii1998]. Prompted to close this relevant methodological gap, in this paper, we present the first computer implementation of the LRTDDFT for NC magnets.
We are motivated by the recent developments in the spintronics and magnonics where an enormous practical potential has been ascertained to the members of this material family [Cheong2024, Rimmler2025]. Moreover, non-collinearity often accompanies other exotic or strongly correlated phases of matter, e.g., superconductivity [6, 2, Bruening2025] and it naturally appears in the description of finite temperature magnetism [Gyorffy1985, Antropov1996]. However, the true attractive force of these magnets stems from their beautiful symmetries [Sandratskii1991] rooted deeply in the modern field theories [Azaria1993, Pradenas2024, Pradenas2025].
@@
A pivotal role in the development of magnonic devices is played by the controllability of the spin-waves’ damping [Chumak2015, Joshi2016, Duine2018]. To begin with, they can scatter considerably on crystal imperfections [Buczek2016, Paischer2021, Paischer2021a, Paischer2024]. Moreover, the NC magnets cannot be fully grasped within the linear spin-wave theory and their magnons undergo intrinsic damping due to two-boson interactions even at absolute zero [Chernyshev2009, Zhitomirsky2013, Park2020] contrary to the case of collinear ferromagnets [Paischer2024]. Finally, in metals, the collective density modes can interact with the single electron-hole pairs which leads to their possibly pronounced Landau decay mentioned above. The mechanism was put forth by Landau to explain the plasmon attenuation [Landau1946] and stems from hybridization of the collective charge and spin density modes with single electron-hole excitations. In the case of collinear magnons, these states involve particles of opposite spins (up holes and down electrons for strong ferromagnets [Paischer2024a]). This simple picture is lost when the transitions stem from the NC band structure and the concept of the Stoner excitation must be revised [Eilmsteiner2026a]. This is achieved natively in the NC LRTDDFT formalism.
We apply our scheme to study the collective spin excitations in an important member of the frustrated NC kagome antiferromagnet (KAFM) family, Mn3Ir [Park2018, Park2020]. Here, we are motivated by the fact that the kagome lattice gives rise to many remarkable properties, including Berry curvature, massive Dirac fermions, quantum Hall effect, topological states, and unconventional superconductivity [Kida2011, Ye2019, Chen2021, Jiang2021, Zhao2021, Guo2022]. From the magnetism point of view, a long overlooked property of the KAFM is their inherently spin-polarized band structure persisting despite the vanishing net magnetization [3, 4], akin to the altermagnetism reported in the collinear magnets [Mazin2022, Smejkal2022, 1, 5, Maznichenko2024a, Sandratskii2025]. It turns out that this NC altermagnetism gives rives to a series of novel magnonic effects, including exotic magnon chiralities and a polarization-dependent Landau damping [Eilmsteiner2026a]. A related phenomenon was reported in collinear compensated ferrimagnets [Odashima2013] but has not been investigated in the NC case whatsoever.
@@
The paper is organized as follows. Sections 2 and 3 expose the underlying formalism and our computer implementation based on the Korringa-Kohn-Rostoker Green’s function method. In section 4, we concentrate on the spin density response of an archetypical KAFM, Mn3Ir. We discuss the magnon dispersion and the chirality-selective Landau damping, and proceed to the real-time and real-space analysis of the spin-wave modes, outlining their prospects in the the terahertz spintronics.
@@ 2.1 Generalized susceptibility
In the linear regime, the quantity yields the charge or magnetization density response $\delta n^{i}{\left({\mathbf{x}},t\right)}$ emerging due to the dynamical perturbing fields $\delta V^{j}{\left({\mathbf{x}}^{\prime},t^{\prime}\right)}$ (magnetic or scalar) coupling to these densities. $\hat{\sigma}^{i}{\left({\mathbf{x}}t\right)}$ are charge ($i=0$) and magnetization density operators ($i=x,y,z$), ${{\left[A,B\right]}}\equiv AB-BA$ is the commutator, and $^{\left\langle\hat{o}\right\rangle}$ is the expectation value of the operator $\hat{o}$ for the system without the perturbation. Throughout all the paper, $\mathbf{x}$} refers to the global position vector, whereas $\mathbf{r}$} is the position within the primitive unit cell. It is assumed that the potentials of the unperturbed system bear no time dependence in the Schrödinger picture. Note that the response function can be used to study any electron density excitations including magnons, excitons, and plasmons.
@@
Here, $\mathbf{r}$} and $\mathbf{r}$$^{\prime}$ belong to the primitive cell of the crystal and $\mathcal{L}$ denotes the set of lattice vectors. (In the equation above, other arguments are suppressed for simplicity.)
@@
allows to analyse the spectra and the spatial forms of the excited density modes. For a specific energy $\omega$, the eigenvectors $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ of $L$ represent the shapes of the natural modes of charge and magnetization density fluctuations at this frequency. The magnitude of the associated eigenvalues, $a_{\lambda}{\left({\mathbf{q}},\omega\right)}$, gives the density of states of these modes, up to a factor $1/\pi$. In other words, large $a_{\lambda}$ signifies the presence of density modes in the system. These modes are real physical objects of the following form
${\mathfrak{Re}{\left[\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}{\mathfrak{e}^{{{\mathfrak{i}}}{\left({\mathbf{q}}\cdot{\mathbf{R}}-\omega t+\phi\right)}}}\right]}},$ (4)
where $\phi$ is an overall phase of the mode. When the frequency dependence of $a_{\lambda}{\left({\mathbf{q}},\omega\right)}$ is narrow, one often identifies such modes as “collective”. Following the fluctuation dissipation theorem, which we briefly outline in Appendix A, the loss matrix has several further useful physical interpretations. On one hand, $\omega a_{\lambda}{\left({\mathbf{q}},\omega\right)}/2$ yields the power absorbed by the system from the driving external fields of the shape $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ coupling to the densities. On the other, it can be used to compute the expectation values of the time correlations ${\mathfrak{C}}^{ij}{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}},\omega\right)}$ between density operators in the unperturbed state, i.e., among others, the spontaneous fluctuations of the system density.
Furthermore, the magnetization density loss matrix evaluated in momentum space is proportional to the cross-section of the inelastic neutron scattering off spin-density excitation [VanHove1954]. Last but not least, the quantity is related to the (spin-polarized) electron energy loss spectroscopy ((SP)EELS) signal [Kirschner1985, Vollmer2003, Costa2006, Santos2018]. In the latter case, however, contrary to the neutron scattering, the loss energy dependence of the cross-section involves intricate details of the Coulomb interaction between the electrons of the beam and the probe [Vignale1985, Plihal1998, Hong2000a, Paischer2024b].
@@ 2.2 Linear response time dependent density function theory formulation
The evaluation of the dynamic susceptibilities of an interacting electron gas is as involved as the solution of the many-body problem itself [Abrikosov1975]. Consequently, one must resort to approximations. A successful ab initio scheme can be derived within the LRTDDFT [Gross1985, Qian2002]. It often involves two steps. First, the so called Kohn-Sham susceptibility, $\chi_{\mathrm{KS}}$ is constructed describing the response of the formally non-interacting Kohn-Sham (KS) system to an external perturbation. Second, the self-consistent dynamical change in the exchange-correlation potential arising due to the densities induced by the external field. This change is incorporated upon solving the so called susceptibility Dyson equation. There are different complementary approaches to implement this scheme of which some do not explicitly break it into these two distinct phases [Savrasov1998, Binci2025]. The magnetic response function can be also found within the many-body perturbation scheme [Aryasetiawan1999, Karlsson2000, Mueller2016, Nabok2021] and using empirical Hubbard-type Hamiltonians [Cooke1980, Gumbs1980, Tang1998, Muniz2002, Bonetti2022].
@@
(For simplicity the arguments of the Green’s functions as well as corresponding sums and integrals are suppressed.) The ground state potential $V$ includes static external fields as well as the ground state Hartree and exchange-correlation contributions. In this case, it does not feature the dependence on the complex energy $z$ and is local. This is the form used to describe the materials considered in this paper.
Additionally, in many cases, it is necessary to include terms originating from the spin-orbit coupling (SOC) [Paischer2024]. Furthermore, the frequency dependence arises when typical electronic self-energy terms are incorporated, as it is the case when the disorder is described in the coherent potential approximation (CPA) [Buczek2016, Paischer2021, Paischer2021a] or many-body effects like magnon-electron interaction are included [Paischer2023, Usachov2024, Paischer2024a]. In the latter case, the self-energy $V$ can additionally become non-local.
@@
${f_{T}}({\epsilon}_{j})$ stands for the Fermi-Dirac distribution function. A small finite temperature is introduced to slightly smear the Fermi level allowing for better convergence of the Brillouin zone integrals. In this work, this temperature is negligible on the electronic energy scales and the electron gas is regarded to be described at absolute zero. $0^{+}$} denotes the small positive energy limit. Einstein summation convention is assumed for the spin indices. There is no difference raised and lowered density index $i$. The advantage of this approach is the possibility of deforming the complex plane integration contours such that the evaluation of $S$ takes place far from the singularities of the KS Green’s function allowing for its more rapid numerical convergence [Buczek2011a]. Furthermore, the evaluation of the susceptibility on the nearly real axis yields a more stable analytic continuation to the real frequencies and magnon line fitting schemes than the evaluation at Matsubara poles [Staunton1999, Staunton2000, Thakor2003] although it is computationally more expensive.
@@ 2.3 Exchange-correlation kernel
The exact form of the $K_{\mathrm{xc}}$ is, in general, unknown, especially for NC systems. There is a steady progress in constructing NC spin density functionals [Sharma2007, Kurth2009, Eich2010, Eich2013, Eich2013a, Pittalis2017, Eich2018], but their their inclusion in practical LRTDDFT schemes is still a work in progress. In this work, we adopt the adiabatic local spin density approximation (ALSDA) yielding $K_{\mathrm{xc}}$ in the following form
@@
For NC systems, this kernel is most easily written in a local coordinate system with its $z$-axis oriented along the magnetization direction [vonBarth1972, Katsnelson2004]. In this case, its space-diagonal part reads
$K_{\mathrm{xc}}^{\textrm{loc}}=\begin{pmatrix}{\frac{\partial V_{\mathrm{xc}}}{\partial n}}&0&0&{\frac{\partial V_{\mathrm{xc}}}{\partial m}}\\ 0&K^{\perp}&0&0\\ 0&0&K^{\perp}&0\\ -{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial n}}&0&0&-{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial m}}\\ \end{pmatrix}$ (12)
@@
$K^{\perp}{\left({\mathbf{x}}\right)}$ $=-{\mu_{\textrm{B}}}\frac{B^{\mathrm{GS}}_{\gamma}{\left({\mathbf{x}}\right)}}{m^{\mathrm{GS}}_{\gamma}{\left({\mathbf{x}}\right)}}.$ (14)
$K^{\perp}$ is found simply as the ratio of the magnitudes of KS magnetic field and magnetization density. We recall that these two latter vectors are parallel everywhere in the LSDA. Furthermore, it is immediately evident that $K^{\perp}$ is isotropic in the subspace of the transverse magnetization direction. This form has a clear physical interpretation. Once a infinitesimal magnetization is induced in the direction perpendicular to ${\mathbf{m}}^{\mathrm{GS}}{\left({\mathbf{x}}\right)}$ the local ${\mathbf{B}}_{\mathrm{xc}}{\left({\mathbf{x}}\right)}$ rotates locally without changing its magnitude.
@@
2.4 The emergence of the Goldstone modes in noncollinear magnets
@@
While the emergence of the three Goldstone bosons in NC triangular ferromagnets can be be deduced from the general symmetry arguments outlined above, it is instructive to inquire how it manifests itself in the LRTDDFT calculations. Here, the correct long-wavelength behaviour, i.e., the presence of a mode or modes with vanishing energy for $q=0$, can be seen as a sum rule arising from the interplay between the exchange correlation kernel $K_{\mathrm{xc}}$ and the static uniform magnetic susceptibility [Lounis2010, Lounis2011, Buczek2011a, Mueller2016]. We name this interplay Goldstone sum rule and it can be stated as follows. The “Dyson denominator” of Eq. (2.2) for $q=0$ and $\omega=0$ (Coulomb interaction, $v_{\textrm{C}}$, plays no role in the formation of the magnetic NGB with anisotropy neglected)
@@
must feature zero eigenvalues. Their multiplicity, $N_{r}$, is equal to the number of linearly independent rigid rotations of the magnetic order parameter which, with spin-orbit coupling neglected, cost no energy. We refer to them as zero-energy magnetic deformations. They are the eigenvectors of the Dyson denominator with the eigenvalue 0. In a NC magnets $N_{r}=3$ while $N_{r}=2$ holds for the collinear case, i.e., it is directly given by the number of broken symmetry generators. However, in the light of the preceding discussion, we note that $N_{r}$ is not necessarily equal to the number of Goldstone modes of a system.
Let us now turn to our specific case of the ALSDA. Apart from having no frequency dependence, this kernel, cf. Sec. 2.3, is $q$-independent and finite at $q=0$. The latter property tends to be a serious drawback in the charge-charge response channel, as it precludes the correct TDDFT description of excitonic effects in semiconductors [Onida2002, Turkowski2009]. On the other hand, it is the correct behaviour in the long wave-length limit for magnetization density channels, yielding correctly the magnon Goldstone mode which would not arise otherwise.
@@
The first is of systematic character and arises often in the many-body perturbation methods [Mueller2016, Nabok2021] where the ground state found self-consistently using an approach (typically LSDA to the ground state functional) different from the one for the kernel (corresponding to the screened Coulomb interaction $W$). Once $W$ were used to self-consistently find the magnetic ground state the deficiency would disappear. Let us briefly remark that the currently available many-body perturbation schemes for computing magnetic excitations are practically equivalent to the LRTDDFT in the ALSDA. They take advantage of the LSDA for the electronic ground state, local and static approximation to the screened Coulomb interaction $W$, as well as the choice of Feynman diagrams yielding the random phase approximation-like equation for the dynamic susceptibility resembling formally the equation (2.2) of the LRTDDFT.